Please Answer all parts

NEedUS Pharmaceuticals is exploring options for shipping a $9.5 million dollar machine to its main plant. Holding cost is high for this type of machinery because of opportunity costs. The machine is not available for production while in transit resulting in lost revenue from a highly profitable drug. The annual holding cost is 40%. Options are: A, 40 day transit costing 95,000, B. 25 day transit costing 160,000, and C, 15 day transit costing 280,000.

a. What is the total cost for shipping option A?

b. What is the total cost for shipping option B?

c. What reduction in direct transit cost is required to make shipping option C equal to the lowest cost option?
Answer #1

Holding cost per day = 9,500,000 X 40% / 365 = 10,410.96

a

Total cost = Shipping cost + holding cost = 95,000 + 10,410.96 X 40 = 511,438.4

b

Total cost = 160,000 + 10,410.96 X 25 = 420,274

c

Reduction c total cost = 280,000 + 10,410.96 X 15 = 436,164.4

Reduction required to make it lower cost than option b = 436,164.4 - 420,274 = 15,890.4

Reduction required is15,890.4